Two particles start together from a point O & slide down straight smooth wires inclined at 30° & 60° to the vertical  plane & on the same side of vertical through . the relative acceleration of second with respect to first will be in magnitude:


a) g/2


b) √3g/2


c) g/√3


d) g


 


 

Grade:12

1 Answers

Febi jobi
5 years ago
For first figure acceleration is gsin30°=g/2 .                For second figure acceleration is gSin30°=√3g/2 .  Fix our frame on  first figureand it is considered to be rest .in figure 2 draw the acceleration of first figure which is equal and opposite in direction  split it into its rectangular conponent and that of second figures acceleration. Horizontal acceleration=0.vertical acceleration=√3g/2cos30°-g/2cos60°=g/2